OpenStudy (anonymous):

how to Solve x-Intercept Form of 3x+2y=6?

OpenStudy (anonymous):

3x = -2y +6 Dividing through by 3, x = (-2/3)y + 2 So the x-intercept is 2

OpenStudy (anonymous):

ok, thanks :DD hepls alot to my homework!

OpenStudy (anonymous):

3x + 2y =6 divide by 6 on both side x/2 + y/3 =1 actually you have to make x/a + y/b =1 form then "a" is your x intercept and "b" is your y intercept. so.. 2 is x -intercept and 3 is y intercept

OpenStudy (anonymous):

thanks!! :DD

OpenStudy (aroub):

if u want to find the x-intercept u have to make the y=0

OpenStudy (anonymous):

ya ...aroup has also given you a good technique....

OpenStudy (aroub):

yes and easier :)

OpenStudy (anonymous):

yeah, Thanks guys! :DDDD

OpenStudy (anonymous):

mga bobo naman kse kayu eh

OpenStudy (anonymous):

mga bobo naman kayu dali lng naman ng MATH!! easy sakin yan eh

OpenStudy (anonymous):

solve please 5+2+3=151022 9+2+4=183652 8+6+3=482466 ...5+4+5=202541 then; answer 7+2+5=? answer pls.